OpenStudy (anonymous):

Solve the equation. 5x – 5 = 4x – 8

OpenStudy (anonymous):

again put the x on one side and the numbers on the other so add 5 and subtract 4x to get 5x-4x=-8+5 x=-3

OpenStudy (anonymous):

http://www.myalgebra.com/default.aspx

OpenStudy (anonymous):

dont use programs like that, you will not learn anything!!!